# TA-Cleaner: A Fine-grained Text Alignment Backdoor Defense Strategy for Multimodal Contrastive Learning

### v2

@@ Abstract
Pre-trained large models for multimodal contrastive learning, such as CLIP, have been widely recognized in the industry as highly susceptible to data-poisoned backdoor attacks. This poses significant risks to downstream model training. In response to such potential threats, finetuning offers a simpler and more efficient defense choice compared to retraining large models with augmented data. In the supervised learning domain, fine-tuning defense strategies can achieve excellent defense performance. However, in the unsupervised and semi-supervised domain, we find that when CLIP faces some complex attack techniques, the existing fine-tuning defense strategy, CleanCLIP, has some limitations on defense performance. The synonym substitution of its text-augmentation is insufficient to enhance the text feature space. To compensate for this weakness, we improve it by proposing a fine-grained Text Alignment Cleaner (TA-Cleaner) to cut off feature connections of backdoor triggers. We randomly select a few samples for positive and negative subtext generation at each epoch of CleanCLIP, and align the subtexts to the images to strengthen the text self-supervision. We evaluate the effectiveness of our TA-Cleaner against six attack algorithms and conduct comprehensive zero-shot classification tests on ImageNet1K. Our experimental results demonstrate that TA-Cleaner achieves state-of-the-art defensiveness among finetuning-based defense techniques. Even when faced with the novel attack technique BadCLIP, our TA-Cleaner outperforms CleanCLIP by reducing the ASR of Top-1 and Top-10 by 52.02% and 63.88%, respectively.
@@ 1 Introduction
In the field of artificial intelligence, contrastive learning serves as a powerful learning paradigm aimed at comparing different representations of data, thereby bringing similar samples closer together in the embedding space while pushing dissimilar samples further apart Chen et al. (2020); Khosla et al. (2020); Gutmann and Hyvärinen (2010). This learning method has achieved significant success in many domains, particularly in natural language processing Gao et al. (2021); Chen et al. (2022) and computer vision Bi et al. (2022); Park et al. (2020). In addition to its application in single-modal data, recent works have extended contrastive learning to multimodal data Zhang et al. (2023); Singh et al. (2023); Yang et al. (2022), training on a vast scale of image-text pairs from the web to achieve joint feature representation and matching between images and text. Pre-trained large models, like CLIP Radford et al. (2021b), ALIGN Chen et al. (2021b), and BASIC Chen et al. (2021a), have learned universal representations from large-scale unlabeled data and performed exceptionally well even without task-specific data, as demonstrated by their impressive zero-shot classification performance on ImageNet Deng et al. (2009). In addition to its foundational applications, contrastive learning frameworks have also been adapted to enhance model interpretability Chen et al. (2024), bolster security measures, particularly against adversarial attacks Liang et al. (2021, 2020); Wei et al. (2018); Liang et al. (2022c, a); Wang et al. (2023); Liu et al. (2023a); He et al. (2023b); Liu et al. (2023c); He et al. (2023a); Muxue et al. ; Lou et al. (2024); Kong et al. (2024); Ma et al. (2021, 2022, 2024), and protect privacy Chen et al. (2023); Liang et al. (2022b); enhancing face obfuscation guided by semantic-aware attribution maps (2023); Guo et al. (2023); Dong et al. (2023). These studies not only broaden the scope of contrastive learning applications but also contribute to increasing the robustness and trustworthiness of AI systems in practical deployments. By fine-tuning these models on specific tasks with a small amount of labeled training samples, high-performance vertical domain applications can be realized quickly.
While multimodal contrastive learning (MCL) exhibits powerful feature alignment capabilities, recent studies have also indicated vulnerabilities of such pre-trained large models to data-poisoning backdoor attacks Gao et al. (2020); Jia et al. (2022); Saha et al. (2022); Carlini and Terzis (2021); Li et al. (2023); Liang et al. (2023); Liu et al. (2023b, d); Liang et al. (2024a, b); Zhang et al. (2024); Zhu et al. (2024a); Liang et al. (2024c). By injecting a minimal amount of poisoned samples during the training data collection phase, attackers can introduce biases in the feature learning of the model easily, thereby affecting its feature alignment performance and downstream tasks Liang et al. (2024d). In detail, attackers slightly modify images (embedding various forms of triggers) and the corresponding texts to quickly capture the feature linkage between the triggers and certain specified text during model training, disrupting correct image-text feature matching. This leads to malicious outputs guided by triggered images during subsequent zero-shot inference stages and also poses a series of potential risks for downstream model finetuning.
(a) Synonyms-Synonyms
(b) Random-Random
(c) Synonyms-Random
Figure 1: The t-SNE visualization of feature embeddings between synonym and random word groups.
To our knowledge, faced with a vast training set of unknown security levels, the defense training against models like CLIP focuses on using different kinds of data augmentation strategies to disrupt potential triggers Bansal et al. (2023); Yang et al. (2024a, 2023a). However, the complete training operations are undoubtedly computationally expensive for both attackers and defenders. Attackers may not only poison the original data but also use dirty data to contaminate pre-trained models. The poisoned model is then posted on the Internet or delivered to users Therefore, another defensive approach worth considering is to utilize a clean, small-scale dataset to fine-tune pre-trained large models, thereby reducing their potential backdoor risks. A classic defense method against the CLIP model, CleanCLIP Bansal et al. (2023), discovered that poisoned images and clean images can cluster into two sets in the feature space, and poisoned samples establish additional feature correlations with target attack text. They close the distance between benign and poisoned sets by separately performing self-supervised learning on images and text, thus disrupting the "shortcut" established by backdoor attacks in the multimodal feature space.
However, in the supervised learning domain, for unimodal data, fine-tuning strategies can achieve extremely excellent defense performance, e.g., DBD Huang et al. (2022) and NAD Li et al. (2021a). So in the semi/unsupervised domain, for MCL, can fine-tuning achieve the same degree of defense performance? We have noticed that (1) when using CleanCLIP as a fine-tuning defense against different attack methods, even though the Top-1 ASR can be reduced to different degrees in the zero-shot classification test, the performance of Top-10 ASR still has a large space of decrease. (2) when facing carefully designed triggers optimized in the feature space by attackers, such as BadCLIP Liang et al. (2023), CleanCLIP shows its powerlessness (Top-10 ASR: 98.81%→89.60%). Although CleanCLIP is a classic and superior multimodal defense technique, are there any optimizations that can be made to improve its performance?
We explored the unimodal augmentation technique used by CleanCLIP iiihttps://github.com/nishadsinghi/CleanCLIP in self-supervised learning to explore the factors limiting its performance. In CleanCLIP’s defense scheme, although there are text self-supervised learning, we found that synonym replacement, one of the text-augmentation processes, may not update the text feature distribution very well. As shown in Figure 1, we used t-SNE Van der Maaten and Hinton (2008) technology to visualize the feature embeddings of synonym groups and random word groups in the CLIP pre-trained model (Vit-B/32). It was found that synonyms have a certain clustering effect in the feature space. The kernel estimated densities (KDE) Terrell and Scott (1992) of the two types of data are 0.0476 (synonym group) and 0.0381 (random word group), respectively. This clustering effect makes the text augmentation of CleanCLIP insufficient to timely dissociate the features of the target attack text when the carefully designed trigger approaches, leading to a decrease in defensive performance.
To mitigate the weakness of CleanCLIP against triggers optimized in the feature space, in this paper, we propose a fine-grained sub-text alignment scheme. By randomly selecting original image-text pairs, generating corresponding positive and negative sub-texts, and then performing fine-grained contrastive learning between image and sub-texts. Our approach not only preserves the clean inference capability of the original pre-trained model but also enhances its defensive performance. Our contributions can be summarized as follows:
  • We find that text augmentation through synonym replacement has minimal impact on updating text feature embeddings, which is insufficient in the feature space to resist and dissociate from carefully designed optimized triggers.
  • Building on CleanCLIP, we propose TA-Cleaner, a positive/negative sub-text augmentation strategy. By randomly selecting a small number of samples and extracting positive and negative subtexts in each finetune epoch, negative sub-samples disrupt backdoor features at a fine-grained level, while positive sub-samples protect the correct feature alignment from being disturbed by defense mechanisms.
  • We evaluate the performance of our proposed method using zero-shot classification tasks on ImageNet1K. Experimental results demonstrate that even when facing the latest attack techniques, our defense strategy can quickly reduce the ASR and effectively protect the original model’s BA from being compromised.
@@ 2 Related work and preliminaries
Contrastive Language-Image Pre-Training (CLIP) CLIP Radford et al. (2021b), released by OpenAI, stands as a prominent representative of MCL. Inspired by mapping images and texts into a shared feature embedding space $\mathbb{R}^{d}$, CLIP enables the model to understand the semantic relationship between them. CLIP involves two encoders: an image encoder $f_{I}:I\rightarrow\mathbb{R}^{d}$ and a text encoder $f_{T}:T\rightarrow\mathbb{R}^{d}$, which transform the image and text data into representations of dimension $d$. The model is pre-trained through contrastive learning, leveraging vast amounts of internet image-text pairs $\{I_{i},T_{i}\}_{i=1}^{N}$ to learn the associations between images and texts. During training, the CLIP model learns a mapping function that projects images and texts into the same feature space. This is achieved by maximizing the similarity between positive pairs (matching images $I_{i}$ and texts $T_{i}$) while minimizing the similarity between negative pairs (mismatched images and texts). This unsupervised joint learning approach enables the CLIP model to achieve superior performance on various visual and language tasks, including image classification, text caption generation, and image retrieval. The mathematical expression for $loss_{Clip}$ can be found in the Appendix B.
@@
Backdoor attacks Backdoor attacks generally refer to the implantation of specific trigger patterns during the model training process, which enables the model to perform normally under normal conditions but exhibit abnormal behavior under specific conditions, such as when the input contains images with trigger patterns. In the domain of supervised learning, backdoor attacks have garnered significant attention, with notable works including BadNet Gu et al. (2017), Blended Chen et al. (2017), SIG Liu et al. (2020), WaNet Nguyen and Tran (2021), and SSBA Li et al. (2021b)]. Backdoor attacks targeting the CLIP model primarily leverage its capability in learning from multimodal data. Attackers can add image-text pairs containing specific trigger patterns to the training data, allowing the model to learn the association between these trigger patterns and abnormal behaviors. Within the domain of MCL, Carlini and Terzis (2021) pioneered the revelation of its vulnerability to backdoor attacks, demonstrating a successful attack on CLIP, for instance, by poisoning merely $0.01\%$ of the data. Concurrently, Yang et al. (2023c) delved into the impact of attacks from different modalities on MCL. Additionally, research on attacks against self-supervised learning (SSL), a broader category, is also ongoing, exemplified by BadEncoder Jia et al. (2022), GhostEncoder Wang et al. (2024), and distribution-preserving attacks Tao et al. (2023). The details about data-poisoning backdoor attacks on CLIP are shown in Appendix A.
@@
Backdoor Defenses on CLIP To address these threats mentioned above, some researchers have borrowed backdoor defense techniques from supervised learning Zhu et al. (2023, 2024b) to mitigate the backdoor effects in MCL models. Currently, defense techniques for MCL can be categorized into two groups based on whether the defender can access the poisoned dataset: ① defenders can access the entire poisoned dataset Yang et al. (2023b, 2024b); Bansal et al. (2023); ② defenders can only access the poisoned model Bansal et al. (2023); Niu et al. (2024). The former approach, which allows for complete retraining of large models with various data augmentation strategies, can achieve strong defense performance, such as RoCLIP Yang et al. (2024b). However, in reality, the feasibility of attackers manipulating the training set is low, as they cannot guarantee that their carefully crafted poisoned data will be incorporated into large-scale training sets. Therefore, a more realistic attack strategy is to perform low-cost fine-tuning of existing pre-trained large models with dirty data. As a result, defense techniques targeting the fine-tuning phase are necessary, which is the attack-defense scenario addressed in this paper. A representative example of such defenses is CleanCLIP Bansal et al. (2023). Specifically, CleanCLIP introduces a self-supervised loss based on multimodal data augmentation, which fine-tunes a clean dataset to reduce the impact of backdoor models. Their self-supervised loss $loss_{SS}$ and total fine-tuning loss $loss_{CClip}$ can be found in Appendix B.
@@ 3.1 Threat model
Adversary Objective: The primary objective of the adversary is to manipulate CLIP’s textual output representation. By polluting the original dataset, the model can generate malicious adversarial text specified by the adversary for any input image embedded with a trigger. During zero-shot testing, the attack objective manifests as poisoned images will be misclassified as the adversarial category, while other benign images will be correctly classified.
@@
Figure 2: The framework of our TA-Cleaner, including the sub-texts generation and finetuning. There are three ways to generate the negative sub-captions, and we select one of them randomly for every raw caption. We apply the text augmentation on a little part of samples during every finetuning epoch.
3.2 Fine-grained text augmentation
To address the weakness of CleanCLIP in facing strongly poisoned triggers, particularly in the context of lower performance in text-modal self-supervision, we propose a Fine-Grained Text Augmentation strategy based on positive and negative sub-captions, as illustrated in the left part of Figure 2. Our text augmentation strategy consists of two main parts: positive sub-caption generation and negative sub-caption generation. Assuming the image-text dataset used by CLIP finetuning is $\mathcal{D}_{ft}$, we annotate each sample as $(I_{i},T_{i})\in\mathcal{D}_{ft}$, where $I_{i}$ is the image and $T_{i}$ is its associated caption. And we generate $[ST^{i}_{p},ST^{i}_{n}]$ for each $T_{i}$.
Positive sub-caption generation For each textual data $T_{i}$, we decompose it by first extracting the relational verbs involved. We select one of these verbs as the central word for the positive sub-text $ST^{i}_{p}$. If the subject and object have corresponding adjectives, we randomly remove or retain them. With the help of SceneGraphParser iiiiiihttps://github.com/vacancy/SceneGraphParser, we generate the positive sub-caption using the following template, ensuring the correct positioning of subjects and objects.

### v3

# CleanerCLIP: Fine-grained Counterfactual Semantic Augmentation for Backdoor Defense in Contrastive Learning

@@ Abstract
Multimodal contrastive models like CLIP are increasingly vulnerable to data-poisoning backdoor attacks. Existing defense methods primarily target the pretraining phase. However, with the rise of open-source communities, pretrained models are now freely available for download and fine-tuning. These models may carry unknown security risks, posing significant threats to downstream users. This highlights the need for lightweight defense strategies tailored specifically for the fine-tuning stage. Current defenses during fine-tuning include: finetuning with clean data; and using unimodal self-supervised techniques like CleanCLIP, which has represented the state-of-the-art (SOTA). However, these methods rely on strengthening clean feature representations to mitigate attacks, making them ineffective against more stealthy backdoor techniques, such as BadCLIP, which leverage covert toxic features. To overcome this limitation, we propose a finetuning defense mechanism based on fine-grained counterfactual text semantic augmentation. By modifying small portions of text during fine-tuning, our approach disrupts the association between backdoor triggers and target features. We evaluate our method against six attack algorithms and conduct comprehensive zero-shot classification on ImageNet1K. Experimental results demonstrate that our method achieves SOTA performance in fine-tuning defense. Specifically, when facing the novel BadCLIP attack, our method surpasses CleanCLIP, reducing the Attack Success Rate (ASR) by 52.02% in the Top-1 and 63.88% in the Top-10 classifications.
@@ 1 Introduction
Contrastive learning serves as a powerful learning paradigm aimed at comparing different representations of data, thereby bringing similar samples closer together in the embedding space while pushing dissimilar samples further apart [7, 17, 14]. In addition to its application in single-modal data [11, 6, 2, 24], recent works have extended contrastive learning to multimodal data [40, 29, 35], training on a vast scale of image-text pairs from the web to achieve joint feature representation and matching between images and text. Multimodal contrastive pre-trained models, such as CLIP [26], ALIGN [5], and BASIC [4], have learned universal representations from large-scale unlabeled data and performed exceptionally well even without task-specific data, as demonstrated by their impressive zero-shot classification performance on ImageNet [9]. By fine-tuning these models on specific tasks with a small amount of labeled training samples, high-performance vertical domain applications can be realized quickly.
Figure 1: Overview of backdoor attack strategies and defenses. (a) Traditional backdoor attacks form pseudo-semantic clusters by linking visual triggers to specific texts. (b) BadCLIP avoids detection by directly targeting true feature regions without creating pseudo-clusters. (c) CleanCLIP disrupts pseudo-clusters using self-supervised learning. (d) CleanerCLIP enhances defense by generating fine-grained counterfactual subtexts, breaking the semantic link between the trigger and target.
However, recent research has revealed that these models are vulnerable to data-poisoning backdoor attacks [12, 16, 27, 3, 19], which can compromise their integrity and reliability. In a backdoor attack, an adversary embeds a trigger into the model, allowing it to misclassify inputs in specific, often harmful ways. This vulnerability poses a serious concern, particularly as these models are increasingly deployed in real-world applications. Existing defense methods primarily target the pretraining phase, aiming to mitigate risks before models are fine-tuned for specific tasks. Notable approaches include CleanCLIP [1] and RoCLIP [37], which focus on enhancing the model’s robustness against backdoor attacks during this initial stage. However, the rise of open-source communities has facilitated the widespread availability of pre-trained models, many of which may harbor unknown security risks. Users often download and fine-tune these models for personalized applications, inadvertently exposing themselves to potential threats. This scenario underscores the critical need for lightweight defense strategies that can be applied during the fine-tuning stage.
Current fine-tuning defenses typically involve two which primarily rely on reinforcing clean feature representations: 1) FT: directly fine-tuning with clean samples. 2) CleanCLIP: employing unimodal self-supervised learning, which can also be adapted for the fine-tuning phase and has achieved SOTA performance. While effective against some known threats, these two approaches become inadequate when facing more covert backdoor techniques, such as BadCLIP [21]. BadCLIP can exploit hidden toxic features within the clean feature space, evading detection by bypassing the self-supervised defenses of methods like CleanCLIP.
To clarify the motivation behind our method and its effectiveness, we explore the landscape of backdoor attacks. Traditional backdoor attacks create new feature clusters in the feature space by linking visual triggers to specific texts, thereby assigning new pseudo-semantics to the target text. While these attacks can be effective, they leave distinct traces of pseudo-clusters, making detection and defense more manageable, as shown in Figure 1(a). CleanCLIP further addresses this vulnerability by incorporating a vision-language self-supervised learning module to disrupt these pseudo-semantic clusters, as shown in Figure 1(c). While effective against some known threats, CleanCLIP performs inadequately when facing more covert backdoor techniques, such as BadCLIP [21], which exploits hidden toxic features to evade detection. Instead of generating new pseudo-clusters, BadCLIP precisely identifies the true feature regions of the target text and adjusts the image trigger to approach these regions, successfully evading the self-supervised enhancements, as shown in Figure 1(b). This limitation underscores the necessity for a more robust defense mechanism.
To address this novel challenge, we introduce CleanerCLIP, an innovative strategy that utilizes fine-grained counterfactual semantic augmentation to disrupt the potential semantic link between the trigger and the target output, as illustrated in Figure 1(d). In contrast to CleanCLIP, which primarily focuses on disrupting pseudo-semantic clusters, our approach also generates negative and positive subtexts for a small subset of the clean fine-tuning data. Negative subtexts are created by randomly replacing components of the text’s semantics. This random alteration disrupts the semantic binding exploited in potential backdoor attacks, reducing the stability and success rate of the trigger. Meanwhile, positive subtexts preserve the essential semantic features of the original target text, ensuring that the model can accurately process clean data. This dual augmentation process not only lowers the success rate of backdoor attacks but also enhances the overall robustness of the model during fine-tuning.
Our contributions can be summarized as follows:
  • We identify the limitations of current backdoor defense methods during fine-tuning, particularly their failure to address covert attacks like BadCLIP, highlighting the need for more robust defenses.
  • We propose CleanerCLIP, a lightweight defense strategy that employs fine-grained counterfactual semantic augmentation to disrupt the trigger-target connection and enhance robustness.
  • Our method is tested against six attack techniques, and experimental results show that CleanerCLIP significantly reduces attack success rates while maintaining the model’s benign accuracy and usability.
@@ 2 Related work and preliminaries
Contrastive Language-Image Pre-Training CLIP [26], released by OpenAI, stands as a prominent representative of MCL. Inspired by mapping images and texts into a shared feature embedding space $\mathbb{R}^{d}$, CLIP enables the model to understand the semantic relationship between them. CLIP involves two encoders: an image encoder $f_{I}:I\rightarrow\mathbb{R}^{d}$ and a text encoder $f_{T}:T\rightarrow\mathbb{R}^{d}$, which transform the image and text data into representations of dimension $d$. The model is pre-trained through contrastive learning, leveraging vast amounts of internet image-text pairs $\{I_{i},T_{i}\}_{i=1}^{N}$ to learn the associations between images and texts. During training, the CLIP model learns a mapping function that projects images and texts into the same feature space. This is achieved by maximizing the similarity between positive pairs (matching images $I_{i}$ and texts $T_{i}$) while minimizing the similarity between negative pairs (mismatched images and texts). This unsupervised joint learning approach enables the CLIP model to achieve superior performance on various visual and language tasks, including image classification, text caption generation, and image retrieval. The mathematical expression for $loss_{Clip}$ can be found in Sec 7 in the supplementary material.
@@
Backdoor attacks Backdoor attacks generally refer to the implantation of specific trigger patterns during the model training process, which enables the model to perform normally under normal conditions but exhibit abnormal behavior under specific conditions, such as when the input contains images with trigger patterns. In the domain of supervised learning, backdoor attacks have garnered significant attention, with notable works including BadNet [13], Blended [8], SIG [22], WaNet [23], and SSBA [20]]. Backdoor attacks targeting the CLIP model primarily leverage its capability to learn from multimodal data. Attackers can add image-text pairs containing specific trigger patterns to the training data, allowing the model to learn the association between these trigger patterns and abnormal behaviors. Within the domain of MCL, [3] pioneered the revelation of its vulnerability to backdoor attacks, demonstrating a successful attack on CLIP, for instance, by poisoning merely $0.01\%$ of the data. Concurrently, [39] delved into the impact of attacks from different modalities on MCL. Additionally, research on attacks against self-supervised learning (SSL), a broader category, is also ongoing, exemplified by BadEncoder [16], GhostEncoder [32], and distribution-preserving attacks [31]. The details about data-poisoning backdoor attacks on CLIP are shown in Sec 6 in the supplementary material.
@@
Backdoor Defenses on CLIP To address these threats mentioned above, some researchers have borrowed backdoor defense techniques from supervised learning [41, 42] to mitigate the backdoor effects in MCL models. Currently, defense techniques for MCL can be categorized into two groups based on whether the defender can access the poisoned dataset: ① defenders can access the entire poisoned dataset [36, 38, 1]; ② defenders can only access the poisoned model [1]. The former approach, which allows for complete retraining of large models with various data augmentation strategies, can achieve strong defense performance, such as RoCLIP [38]. However, in reality, the feasibility of attackers manipulating the training set is low, as they cannot guarantee that their carefully crafted poisoned data will be incorporated into large-scale training sets. Therefore, a more realistic attack strategy is to perform low-cost fine-tuning of existing pre-trained large models with dirty data. As a result, defense techniques targeting the fine-tuning phase are necessary, which is the attack-defense scenario addressed in this paper. A representative example of such defenses is CleanCLIP [1]. Specifically, CleanCLIP introduces a self-supervised loss based on multimodal data augmentation, which fine-tunes a clean dataset to reduce the impact of backdoor models. Their self-supervised loss $loss_{SS}$ and total fine-tuning loss $loss_{CClip}$ can be found in Sec 7 in the supplementary material.
@@ 3.1 Threat model
Figure 2: The framework of our CleanerCLIP, illustrating the process of factual(positive) and counterfactual(negative) sub-text generation and fine-tuning. For each raw caption, one of three counterfactual generation strategies is randomly applied. Text augmentation is selectively performed on a small portion of samples during each fine-tuning epoch to ensure minimal computational overhead.
Adversary Objective: By polluting the original dataset, the model can generate malicious adversarial text specified by the adversary for any input image embedded with a trigger. During zero-shot testing, the attack objective manifests as poisoned images will be misclassified as the adversarial category, while other benign images will be correctly classified.
@@
Defender Capability: The defender has access to the model structure and a clean fine-tuning dataset, and can retrain the model to mitigate the attack. However, the defender cannot access the model’s parameters
3.2 Counterfactual Text Augmentation
To address the inadequacies of existing defenses like CleanCLIP against covert backdoor attacks, particularly those employing stealthy trigger features, we propose CleanerCLIP, a fine-grained counterfactual text augmentation strategy. Our approach recognizes the necessity of both preserving clean sample characteristics and disrupting the malicious semantic links exploited by potential backdoor attacks. As illustrated in the left part of Figure 2, our method consists of two main parts: (1) Factual positive sub-caption generation, which ensures the integrity of the original clean data; (2) Counterfactual negative sub-caption generation, which actively undermines the stealthy backdoor triggers. For convenience, we will refer to positive sub-captions as factual sub-captions and negative sub-captions as counterfactual sub-captions in the following descriptions. Assuming the image-text dataset used by CLIP finetuning is $\mathcal{D}_{ft}$, we annotate each sample as $(I_{i},T_{i})\in\mathcal{D}_{ft}$, where $I_{i}$ is the image and $T_{i}$ is its associated caption. And we generate $[ST^{i}_{p},ST^{i}_{n}]$ for each $T_{i}$, representing the positive sub-caption and negative sub-caption respectively.
Factual: Positive sub-caption generation For each text sample $T_{i}$, we decompose it by first identifying its core semantic components, focusing on relational verbs and key nouns. To generate the positive sub-caption $ST^{i}_{p}$, we retain the primary relational verb and key nouns while selectively simplifying or omitting adjectives, ensuring that the core semantic meaning of the original text remains intact. Using SceneGraphParser  [^1] https://github.com/vacancy/SceneGraphParser, we organize the positive sub-caption in the following template, preserving the correct relationships between subjects and objects: